he average airbag is 56 liters. If we're using our baking soda and vinegar airbag, how many grams of sodium bicarbonate are needed to fully inflate the bag? Assume that one mole of gas occupies a volume of 24.5 L at room temperature and 1 atm.
1
Expert's answer
2020-01-28T06:15:46-0500

Sodium bicarbonate, NaHCO3 molar mass=84g

If 56L fully inflate the airbag, then 56000cm3 of gas full inflate the airbag.

And 56/24.5 =2.286moles of gas fill the airbag. Therefore;

Mass of NaHCO3 needed=2.286*84= 192.024g